Photofragment Spectroscopy of HI Photodissociation.

1. Doppler Spectroscopy of the Hydrogen Fragment

\underline{Denis J. Gendron}, John W. Hepburn, CMBLC, Dept. of Chemistry, University of Waterloo, Ontario, CANADA.

The photodissociation of HI has been investigated at various photolysis wavelengths in the first continuum. The dynamic of the process was probed by Doppler spectroscopy at the Lyman α transition of the H atom products. The β-parameters and the branching ratio of the two product channels have been measured at 8 photolysis wavelengths across the first continuum of HI.